OpenStudy (anonymous):

The daily profit made by a game manufacturer is modelled by the equation P=-10(x-60)(x-15), Where P dollars is the profit and x dollars is the price of each game. a) Determine the x intercepts of this equation. b) What do the x intercepts represent in this situation ? Explain. The daily profit made by a game manufacturer is modelled by the equation P=-10(x-60)(x-15), Where P dollars is the profit and x dollars is the price of each game. a) Determine the x intercepts of this equation. b) What do the x intercepts represent in this situation ? Explain. @Mathematics

OpenStudy (anonymous):

Set each factor equal to 0: (x-60) = 0 x = 60 (x-15) = 0 x = 15 So the x-intercepts are x=15, 60 These stand for the prices that if a store sold the games at, they would not make any money.

OpenStudy (anonymous):

so they would be the Max and Min ?

OpenStudy (anonymous):

Not really. They are the prices that if they sell games at that price then they will not make any money and they will not lose any money.

OpenStudy (anonymous):

oh ok, Thanks!